OpenStudy (anonymous):

Which reforms to make the government more democratic were supported by the Progressive movement? A. direct primaries, initiative, and referendum B. term limits, limits on campaign spending, and women's voting C. women's suffrage, a progressive income tax, and the youth vote D. the end of the Electoral College, voting rights for Indians, and a volunteer military

OpenStudy (anonymous):

Which was not a change taking place in the United States in the late 1800s? A. Blacks and other minorities were gaining equality. B. People had more leisure time. C. Factories were mass-producing conveniences and luxuries. D. The middle class was growing.

OpenStudy (anonymous):

@texaschic101 can you help please

OpenStudy (anonymous):

can u help me plz i will give u a medal

OpenStudy (anonymous):

19,552,860 is what percent of 316,128,839

OpenStudy (anonymous):

Trevor go look at your posted question I answered there

OpenStudy (texaschic101):

1. A 2. I am not sure about this one....they all seem correct to me.

OpenStudy (texaschic101):

is/of = %/100 19,552,860/316,128,839 = x/100 cross multiply 316,128,839x = 19,552,860(100) 316,128,839x = 1,955,286,000 x = 1,955,286,000/316,128,839 x = 6.185 rounds to 6.19%

OpenStudy (texaschic101):

that might be wrong.....a lot of numbers....I just plugged in the numbers into the formula.